OpenStudy (anonymous):

What is the very general formula for finding the slope of an asymptote in a hyperbola?

OpenStudy (anonymous):

well lets consider \[f(x)=a/(x-p)+p\] \[f \prime (x)=-a/(x-p)^{2}\]

OpenStudy (anonymous):

which asymptote vertical or horizontal

OpenStudy (amistre64):

\[\frac{(x-C_x)^2}{a^2}\pm\frac{(y-C_y)^2}{b^2}=1\] \[A=\pm\sqrt\frac{b^2}{a^2}(x-C_x)+C_y\]maybe